# Edge-Only Universal Adversarial Attacks in Distributed Learning

### v1

@@ 4.2 Adversarial Optimization
5 Experimental Evaluation
After introducing the used settings, the section addresses the following research questions: 1 Are edge-only UAPs capable of propagating a strong misprediction effect to the cloud part, and how do they compare with a full-knowledge UAP formulation [39, 20]?; and 2 How the targeted objective in edge-layers propagates to the unknown cloud part?
Ablations studies were also conducted, aimed at validating the effectiveness of the attacks under different settings.
5.1 Settings
@@
All the experiments were conducted on the ImageNet dataset [14]. We used class-balanced subsets of the ImageNet validation set: 5000 images as a test set for evaluating all attacks, 1000 images for $\mathcal{D}_{o}$, and other 1000 images for $\mathcal{D}_{\text{opt}}$. While $\mathcal{D}_{t}$ was composed of only 50 samples for each selected target class, these also extracted from the validation set. For the attacked models, we selected CNNs architectures commonly used in distributed scenarios, including ResNet50 [11], Wide-ResNet101 [38], MobileNetV2 [31], and VGG16 [33] [^1] Weights loaded from torchvision..
@@
To avoid an overly fine-grained analysis of internal layers, where feature representations between adjacent layers may exhibit minimal variation, we selected four key layers in each network, representing different depths within the backbone of each model. Specifically, these layers are chosen at depths of $1/4$, $2/4$, $3/4$, and $/4$ 4 of the backbone. Each of these layers serves as a potential split point, defining four configurations of the edge and cloud components, $f_{\textit{edge}}$ and $f_{\textit{cloud}}$, corresponding to different edge depths. For clarity, we denote a configuration of the edge component by the depth of the selected key layer that serves as the split point. For example, an edge depth of $1/4$ means that the edge component includes all layers up to the $1/4$ depth of the backbone (e.g., the first residual block in the ResNet architecture). Conversely, an edge depth of $/4$ 4 indicates that the edge part covers the entire backbone of the model.
To simplify the notation, $l_{i}$ is used in the following to denote key layers only. As such, given a depth of the edge part, the set $L_{\text{edge}}=\{l_{1},...,l_{e}\}$ defined in Section 4 is formed by the key layer $l_{e}$ acting as a split point and the preceding key layers. For instance, for depth 2/4, the key layers are $L_{\text{edge}}=\{l_{1},l_{2}\}$, where $l_{1}$ and $l_{2}$ correspond, for example, to the 3rd and 7th layer of the network, respectively.
@@
The binary classifiers $g^{l}$ consist of a simple and lightweight module (avoiding overfitting) that includes: a conv block with kernel size 3, followed by an AvgPool and two fully connected layers with hidden size 128, separated by a ReLU activation. More details about the classifiers and the split points are in the supplementary material.
@@
Regarding the attack settings, we focused on $l_{\infty}$ norm attacks, using $\alpha=2/255$ and the maximum number of epochs for the UAP optimization to $20$, with a batch size of $100$. In the following, we provide comparisons of edge-only UAPs with respect to full-knowledge, state-of-the-art attacks [39]. All experiments were conducted using an A100 GPU with 40GB of memory.
5.2 Untargeted Effectiveness
@@
To address research question 1, Figure 2 shows the effect on the test set accuracy when applying edge-only UAPs with magnitudes of $\epsilon=\frac{10}{255}$ and $\epsilon=\frac{16}{255}$. In particular, we selected five random target classes from ImageNet and then compared edge-only attacks under different depths of the edge part. For comparison, we also included classic, batch-level UAP attacks (both targeted and untargeted) [39] implemented through a cross-entropy loss. Please note that classic UAP assumes full model knowledge, whereas our edge-only attacks requires knowledge of the edge part only. Furthermore, UAP untargeted attacks assume full knowledge of the original labels of the tested samples, which are not available in edge-only attacks by definition (thus limiting their applicability in a targeted setting).
@@
As reasonably expected given the nature of the attack, attacking a shallow edge part, hence basing on limited model knowledge, generally reduces the attack effectiveness (i.e., higher prediction accuracy) compared to cases with access to more layers of the distributed model. This trend is more pronounced for models like ResNet50 and Wide-ResNet101, while different observations arise for MobileNetV2 and VGG16, where edge-only UAPs, even when targeting the initial layers only, achieve similar effectiveness to full-knowledge attacks.
Additionally, analogously as found in previous work[20] that suggests the existence of dominant labels that occupy large regions in the image space, we observed that the choice of target class significantly impacts the generation of strong UAPs. This phenomenon is especially pronounced in edge-only attacks (e.g., see classes ’warthog’ and ’crab’ in Figure 2), where less-dominant classes result in even lower transferability.
@@
depth $/4$ 4
@@
Figure 5: Impact of target edge-only UAP ($\epsilon=10/255$). In (a), we present a TSNE of ResNet50 features, showing 50 samples from the target class (’brain coral’, blue) and 1000 from other classes (red). A new sample (yellow) and its edge-only perturbed versions at different depths are projected. In (b), class prediction frequency is shown across 5000 samples, comparing no attack (green) to edge-only attacks at different depths. The first row shows MobileNetV2 results (target: ’tench’) and the second row shows ResNet50 (target: ’brain coral’).
5.3 Understanding Target Features Across Layers
While the proposed attacks may produce effective mispredictions, we also aim to assess the extent to which they produce cloud outputs aligned with the intended targeted objective 2 and how much steering the edge features towards a target may impact the remaining decision process of the model.
@@ 5.3 Understanding Target Features Across Layers
To directly evaluate whether the target objective is propagated to the cloud output, Figure 3 reports the target success rate (TSR) on the test set (i.e., the percentage of predictions matching the target class) when using edge-only UAPs crafted for various depths of the edge part. From the results, it is clear that (a) low-depth attacks do not induce mispredictions toward the target class. In fact, only attacks applied to the last layers of the backbone achieve a higher target success rate, which further increases when the attack is directly applied to the model logits (‘cls’, assuming full model knowledge).
@@
To gain deeper insights into how much the target features computed in the early layers move the perturbed samples closer to the target class, we plot in Figure 5(a) a projection of the features extracted by ResNet50’s key layers using T-SNE, distinguish between target and non targeted training samples. Then, we projected in the same space a clear test sample (yellow point) alongside its perturbed versions (green points), when using edge-only UAPs computed at different edge depths ($1/4$, $2/4$, $3/4$, and $/4$ 4), as indicated by the labels in the plot. Note that the features extracted at lower depths are less separable, and all the perturbed versions move the original sample in a region closer to the target ones. However, this trend does not hold while moving in the deeper feature space, as shown in Figure 5(a) with depths $3/4$ and $/4$ 4, where early-stage perturbation move the test sample not sufficiently close to target regions. This analysis suggests that (b) finding the correct direction without knowledge of the remaining processing part of the model could be particularly challenging, as perturbations computed in early stages cannot guide the sample toward the final target features subspaces.
@@
Finally, even though the proposed UAPs do not fully propagate the targeted effect to the cloud part, we wanted to understand their output effect. For this purpose, we plotted the five most frequent predictions from the test set in Figure 5(b) with MobileNetV2 (top Fig., target ’tench’) and ResNet50 (bottom Fig., target ’brain coral’). Each bar plot in the figures shows the results of an attack performed at different edge depths, while the leftmost bars in green denote the no-attack baseline. As expected, without perturbations, the distribution of the prediction remains fairly balanced (the most frequent class is predicted only about 10 times), consistently with the high clean accuracy of the model. However, this frequency distribution drastically changes when perturbations are applied (note values on the x-axis), even at early layers (first model depth). Although initial-layer perturbations may not steer the prediction towards the target class, (c) in the cloud part the attack tends to induce a significant shift in features towards a different reference classes. This insight suggests an interesting direction for future investigation, focusing on potential intra-class relationships that could guide the attack towards a target class by optimizing for a different one.
5.4 Ablation Studies
On the Attack Optimization Steps.
@@ 5.4 Ablation Studies
We conducted ablation studies on the proposed attacks to separately assess the steps involved in optimizing edge-only perturbations (Section 4.2). Specifically, in Table 1 we reported the model accuracy under edge-only UAPs obtained under different settings. In particular, we evaluated the impact of combining the gradient of multiple edge layers during optimization ($\sum_{l\in L_{\text{edge}}}$ in Eq. (6)) and the effect of gradient normalization (denominator of the last term of Eq. (6)), referred to as ”MultiL” and ”Norm” in Table 1, respectively. As shown in the table, applying both these approaches results in improved attack effectiveness. The use of multiple layers helps reduce overfitting in the feature separation analysis within deeper layers, while normalization helps maintain balanced gradients across all layers.
@@
Table 1: Ablation studies on the impact of multiple edge layers’ gradients and gradient normalization. Attacks were performed on MobileNetV2 and ResNet50, with $\epsilon=10/255$ and ’tench’ as the target class.
Figure 6: Analysis of the effectiveness of edge-only universal perturbation when considering different $\epsilon$-magnitude on ResNet50 and MobileNetV2.
Variation of the success rate with $\epsilon$.
In Figure 6, we reported the model accuracy when varying the attack magnitude $\epsilon$ for edge-only UAPs across different layers. For this test, we consider the class ’tench’ as the target reference class for both ResNet50 and MobileNetV2. As illustrated in the plot, the effectiveness of the attack varies significantly, while $\epsilon=\frac{16}{255}$ appears to be a reasonable compromise for high effectiveness and low magnitude. Additional results are provided in the supplementary material, while Figure 8 includes visual examples of perturbed images generated from the first layers (edge depth 1/4) of MobileNetV2, targeting mostly high-frequency components, as shown in Figure 9, due to the nature of early network features.
@@
Figure 8: Illustrations of the proposed edge-only attacks (class = ’tench’) on MobileNetv2 using different manigtudes. Predicted classes and confidences are shown on top. The perturbations are obtained considering the edge part with depth 1/4 of the backbone.
(a) depth $1/4$
(b) depth $2/4$
(c) depth $3/4$
(d) depth $/4$ 4
Figure 9: Edge-only UAPs for different edge size of ResNet50 (top) and MobileNetV2 (bottom).
6 Conclusion and Future Work
In this work, we proposed, to the best of our knowledge, the first edge-only adversarial attack aimed at crafting universal adversarial perturbations without any knowledge of the cloud part model or the entire class distribution. This is accomplished by first training simple classifiers to identify target features in edge layers, and then performing adversarial optimization towards these features.

### v2

@@ V-C Untargeted Adversarial Attack
VI Experimental Evaluation
After introducing the experimental setup, this section presents results for the proposed attacks. The analysis investigates how both the untargeted and targeted formulations are able to induce strong misclassification effects in the cloud-side model, and how their performance compares with classic UAPs. Then, for the targeted setting, we conduct analyses to understand how objectives enforced at the edge-layer representations propagate through the unknown cloud part of the network. Finally, ablation studies and additional experiments were conducted, aimed at validating the effectiveness of the attacks and the threat model proposed.
VI-A Experimental Settings
@@
All the experiments were conducted on the ImageNet dataset [26]. We used class-balanced subsets of the ImageNet validation set: 5000 images as a test set for evaluating all attacks, 1000 images for $\mathcal{D}_{o}$, and other 1000 images for $\mathcal{D}_{\text{opt}}$. While $\mathcal{D}_{t}$ was composed of only 50 samples for each selected target class, these also extracted from the validation set. For the attacked models, we selected CNNs architectures commonly used in distributed scenarios: ResNet50 [48], Wide-ResNet101 [49], MobileNetV2 [50], and VGG16 [51] [^2] Weights loaded from torchvision.. Additionally, we provide experimental results with the integration of different datasets for out-of-distribution analysis (CIFAR-10), as well as the use of ViT models to extend the analysis beyond the classic CNNs typically used in distributed settings (Sec. VI-F).
@@
To avoid an overly fine-grained analysis of internal layers, where feature representations between adjacent layers may exhibit minimal variation, we selected four key layers in each network, representing different depths within the backbone of each model. Specifically, these layers are chosen at $4$ depths of the backbone. Each of these layers serves as a potential split point, defining four configurations of the edge and cloud components, $f_{\textit{edge}}$ and $f_{\textit{cloud}}$, corresponding to different edge depths. For clarity, we denote the configuration of the edge component by the depth of the selected key layer that serves as the split point. For example, an edge depth of $1/4$ means that the edge component includes all layers up to the first major block of the backbone (e.g., the first block in the ResNet architecture). Conversely, an edge depth of $4/4$ indicates that the edge part covers the entire backbone of the model, with the cloud part consisting only of the classification head. Details about the specific layers selected for each model in this configuration are available in the appendix and in the provided source code. Note also that, given the edge depth, the set $L_{\text{edge}}=\{l_{1},...,l_{e}\}$ (defined in Sec. V) is formed by the key layer $l_{e}$ acting as a split point and the preceding key layers.
@@
For the targeted attack presented, the binary classifiers $g^{l}$ consist of a simple and lightweight module (avoiding overfitting) that includes: a conv block with kernel size 3, followed by an AvgPool and two fully connected layers with hidden size 128, separated by a ReLU activation. More details about the classifiers are provided in the appendix, while related ablation studies are in Sec..VI-E3.
@@
Regarding the attack settings, we focused on $l_{\infty}$ norm attacks with different epsilon values, using $\alpha=2/255$ and the number of epochs for the UAP optimization to $20$, with a batch size of $100$. All experiments were conducted using an A100 GPU with 40GB of memory.
VI-B Untargeted Effectiveness
@@
To analyze the degradation in accuracy induced by the proposed UAPs in both targeted and untargeted settings, Fig. 3 reports the impact on test-set accuracy when applying edge-only UAPs with perturbation magnitudes of $\epsilon=\frac{10}{255}$ and $\epsilon=\frac{16}{255}$. For the targeted formulation, we randomly selected five target classes from ImageNet and evaluated the effectiveness of edge-only attacks across different depths of the edge portion of the model. For comparison, we also included classical UAP attacks (both targeted and untargeted) based on a cross-entropy loss [32]. It is important to note that classical UAP methods assume full knowledge of the model and labels. For the untargeted analysis, we evaluated our extension of the CosUAP attack [10] in an edge setting, as discussed in Sec. 8.
@@
As reasonably expected given the nature of the attack, attacking a shallow edge part, hence basing on limited model knowledge, generally reduces the attack effectiveness (i.e., higher prediction accuracy) compared to cases with access to more layers of the distributed model. This trend is more pronounced for models like ResNet50 and Wide-ResNet101, while different observations arise for MobileNetV2 and VGG16, where edge-only UAPs, even when targeting the initial layers only, achieve similar effectiveness to full-knowledge attacks. Additionally, in line with prior work [9], which suggests the existence of dominant labels occupying large regions of the image space, we observed that the choice of the target class has a significant impact on the strength of the resulting UAPs. This effect is especially pronounced in the edge-only setting (e.g., see the classes “warthog” and “crab” in Fig. 3), where less dominant classes have lower transferability.
Different considerations arise for untargeted approaches, such as CosUAP-Edge (shown in the plot), which tend to induce a larger drop in accuracy even when applied at the first or second edge configurations, without requiring the selection of a highly transferable class. This demonstrates the benefits of CosUAP-Edge attacks; however, they do not provide any control over the specific target class.
@@
depth $4/4$
depth $1/4$
depth $2/4$
depth $3/4$
depth $4/4$
@@
Figure 6: Impact of the targeted edge-only UAP ($\epsilon=10/255$) in the layers’embedding space with t-SNE . Visualizations of MobileNetV2 (top) and ResNet50 (bottom) features show 50 samples from the target class (“brain coral,” blue) and 1000 samples from other classes (red). A new test sample (yellow) and its edge-only targeted perturbed versions (green), generated at different depths (numbers).
Figure 7: Class prediction frequency is shown across 5000 samples, comparing no attack (green) to edge-only attacks at different depths. The first row shows MobileNetV2 results (target: ’tench’) and the second row shows ResNet50 (target: ’brain coral’).
VI-C Understanding Target Features Across Layers
@@
While the proposed targeted attacks produce effective mispredictions, we also aim to assess the extent to which cloud outputs align with the intended targeted objective and how much steering the edge features towards a target may impact the remaining decision process of the model. To evaluate this, Fig. 4 reports the target success rate (TSR) on the test set (i.e., the percentage of predictions matching the target class) when using edge-only targeted UAPs crafted for various depths of the edge part. From the results, it is clear that (a) low-depth attacks do not induce mispredictions toward the target class. In fact, only attacks applied to the last layers of the backbone achieve a higher target success rate, which further increases when the attack is directly applied to the model logits (‘cls’, assuming full model knowledge).
@@
To gain deeper insights into how much the target features computed in the early layers move the perturbed samples closer to the target class, we plot in Fig. 6 a projection of the features extracted by ResNet50’s key layers using T-SNE, distinguishing between target and non-targeted training samples. Then, we projected in the same space a clear test sample (yellow point) alongside its perturbed versions (green points), when using edge-only UAPs computed at different edge depths ($1/4$, $2/4$, $3/4$, and $4/4$), as indicated by the labels in the plot. Note that the features extracted at lower depths are less separable, and all the perturbed versions move the original sample in a region closer to the target ones. However, this trend does not hold while moving in the deeper feature space, as shown in Fig. 6 with depths $3/4$ and $4/4$, where early-stage perturbation moves the test sample not sufficiently close to target regions. This analysis suggests that (b) finding the correct direction without knowledge of the remaining processing part of the model could be challenging, as perturbations computed in early stages cannot guide the sample toward the final target features subspaces.
Figure 8: Attacked images on ResNet50 with edge-only UAPs at different depths (1/4, 2/4, 3/4, 4/4), $\epsilon=10/255$ and target ’brain coral’.
(a) depth $1/4$
(b) depth $2/4$
(c) depth $3/4$
(d) depth $4/4$
Figure 10: Edge-only UAPs for different edge sizes of ResNet50 (top) and MobileNetV2 (bottom).
@@
Finally, even though the proposed UAPs do not fully propagate the targeted effect to the cloud part, we wanted to understand their output effect. For this purpose, we plotted the five most frequent predictions from the test set in Fig.7 with MobileNetV2 (top Fig., target ’tench’) and ResNet50 (bottom Fig., target ’brain coral’). Each bar plot in the figures shows the results of an attack performed at different edge depths, while the leftmost bars in green denote the no-attack baseline. As expected, without perturbations, the distribution of the prediction remains fairly balanced (the most frequent class is predicted only about 10 times), consistently with the high clean accuracy of the model. However, this frequency distribution drastically changes when perturbations are applied (note values on the x-axis), even at early layers (first model depth). Although initial-layer perturbations may not steer the prediction towards the target class, (c) in the cloud part the attack tends to induce a significant shift in features towards a different reference classes. This insight suggests an interesting direction for future investigation, focusing on potential intra-class relationships that could guide the attack towards a target class by optimizing for a different one.
VI-D Visualization of the Perturbations.
In Fig. 8, we report some visualizations of the attack with $\epsilon=10/255$, showing perturbed input samples alongside their corresponding predictions and confidence scores for ’brain coral’ as the target class, generated from different edge depths.
As shown in the figure, the ability to successfully induce an output prediction consistent with the target class varies depending on the class itself, with some classes being inherently more difficult to reach. Additionally, it is worth noting that perturbations crafted from shallower layers tend to emphasize high-frequency components. This results in more visually perceptible and aggressive perturbations, which is expected, as shallow features are more fine-grained and sensitive to low-level details. This observation is further illustrated in Fig. 10, where we visualize UAPs generated from four different layers in both ResNet50 and MobileNetV2.
VI-E Ablation Studies
VI-E1 Attack Optimization Steps
@@
We conducted ablation studies on the proposed attacks to separately assess the steps involved in optimizing edge-only perturbations (Sec. V-B). Specifically, in Table I we reported the model accuracy under edge-only UAPs obtained under different settings. We evaluated the impact of combining the gradient of multiple edge layers during optimization ($\sum_{l\in L_{\text{edge}}}$ in Eq. (6)) and the effect of gradient normalization (denominator of the last term of Eq. (6)), referred to as ”MultiL” and ”Norm” in Table I, respectively. As shown in the table, applying both these approaches results in improved attack effectiveness. The use of multiple layers helps reduce overfitting in the feature separation analysis within deeper layers, while normalization helps maintain balanced gradients across all layers.
@@
TABLE I: Ablation studies on the impact (in terms of accuracy on ImageNet) of multiple edge layers’ gradients and gradient normalization. Attacks were performed on MobileNetV2 and ResNet50, with $\epsilon=10/255$ and ’tench’ as the target class.
VI-E2 Varying the number of samples
We conducted additional analysis to better understand the variation in attack effectiveness when changing the number of samples used to train the binary classifiers, while keeping the number of optimization iterations fixed. In particular, Fig. 11 shows an analysis where the number of samples in the ”other dataset” (non-target classes) is fixed at 1000, while the number of target-class samples varies. As shown in the plot, results on ResNet50 and MobileNetV2 indicate that, in general, having more target samples can be beneficial, as the classifier learns to generalize better across different classes.
Figure 11: Analysis of the accuracy across different edge-only attacks (edge having 1/4 and 2/4 sizes of the backbone) when varying the number of target samples (top part) and the number of other samples (bottom part) used to learn feature separations. The attack was conducted with target class 0 (’tench’) and $\epsilon=10/255$.
A different trend emerges when varying the number of other-samples (non-target samples), as shown in Fig. 11 (bottom), where the number of target samples is fixed at 50. No notable change was observed with ResNet50, where the results appear approximately stable, suggesting that varying the number of non-target samples does not significantly affect model performance. However, MobileNetV2 shows a different behavior, with fewer non-target samples proving more beneficial for the attack. We believe that intrinsic properties related to the selection of non-target samples could impact specific models, and we hope to investigate possible sample relationships and intriguing aspects in more detail in the future.
VI-E3 On the size of the binary classifiers
We conduct ablation studies to assess the impact of selecting simple or complex binary classifiers for learning feature separation in the edge space. Specifically, in Fig. 12, we illustrate the accuracy reduction observed when performing the proposed attacks at two different edge stages while varying the depth of the binary classifier, ranging from a single convolutional block (#1), as in the standard setting, to four convolutional blocks (#4). Please see the Appendix for additional architecture details. As shown in the bar plots, while increasing the number of convolutional blocks may improve feature separation, this also shifts the binary classifier toward further processing the features rather than preserving the separation learned by the edge part of the model. As a result, the attack’s effectiveness decreases when leveraging the edge model’s gradients. Therefore, to enhance transferability to the cloud part, a simpler configuration proves to be more effective.
Figure 12: Analysis of the complexity of binary classifiers whit ResNet50 on ImageNet for different target classes (343 and 0). The bars compare the number of convolutional blocks used to perform edge-side attacks from different edge depths.
VI-F Additional test
VI-F1 Variation of the success rate with $\epsilon$
In Fig. 13, we reported the model accuracy when varying the attack magnitude $\epsilon$ for edge-only UAPs across different layers. For this test, we considered the class ’tench’ as the target reference class for both ResNet50 and MobileNetV2. As illustrated in the plot, the effectiveness of the attack varies significantly, while $\epsilon=\frac{16}{255}$ appears to be a reasonable compromise for high effectiveness and low magnitude. A clarification of the magnitude of the features of the edge-side attacks is shown in Fig. 15, which includes visual examples of perturbed images generated from the first layers (edge depth 1/4) of MobileNetV2, targeting mostly high-frequency components, as shown in Fig. 10, due to the nature of early network features.
Figure 13: Analysis of the effectiveness of edge-only universal perturbation whit different $\epsilon$-magnitude on ResNet50 and MobileNetV2.
@@ VI-F1 Variation of the success rate with $\epsilon$
Figure 15: Illustrations of the edge-only attacks (class = ’tench’) on MobileNetv2 using different $\epsilon$. Predicted classes and confidences are shown on top. The perturbations are obtained considering the edge part with depth 1/4 of the backbone.
VI-F2 Analysis of the out-of-distribution attack dataset
In the experimental settings studied in the main experimental part, we evaluated the proposed attack using images from the validation set of ImageNet for both the target and other-samples optimization datasets. Although these images were not used to train the tested model, their distribution may share some similarities, such as resolution and source characteristics, given that they originate from the same dataset (despite the vast diversity of ImageNet). To better investigate the impact of using completely different images, we considered in Table II a dataset with a distinct distribution for the other-set samples used to learn feature separation from non-target classes. Specifically, we used the CIFAR-10 test set to form the other-sample set while retaining only the 50 target images from the ImageNet validation set. As shown in Table II, the attack remains effective in reducing accuracy even when using entirely different images in the other-sample set. However, a decrease in effectiveness is observed, likely due to the higher distributional differences between CIFAR-10 images and the high-resolution images learned by the model. Nevertheless, the binary classifier still detects separability from the target samples, leading to effective untargeted attacks on the model’s output. Differences with respect to Fig. 3 for $\epsilon=16/255$ are shown in parentheses in Table II.
Depth-1 Depth-2 Depth-3 Depth-4
MNV2 1.70 (+1.24) 3.12 (+1.86) 4.24 (+3.14) 5.49 (+3.77)
VGG16 10.39 (-2.88) 7.33 (+4.83) 8.31 (+6.41) 8.07 (+6.59)
RN50 21.56 (+2.06) 12.01 (+3.26) 17.02 (+9.16) 16.12 (+9.41)
TABLE II: Evaluation using the $\mathcal{D}_{o}$ set with a completely different distribution (CIFAR-10). The values represent the accuracy on ImageNet, while the values in parentheses indicate the difference compared to the results obtained in the standard experimental settings shown in Figure 3, with $\epsilon$=16/255 and target class 0.
VI-F3 Testing Edge-only UAPs on ViT
We also evaluate the effectiveness of the proposed attacks across different architectures that, while not frequently addressed in the IoT literature, are still relevant for understanding the attack’s impact and exploring its applicability in future scenarios. Specifically, we considered transformer-based models, such as Vision Transformers (ViTs) [52], in both Base (ViT-B) and Small (ViT-S) versions, applying the same analysis as in previous settings across different layers of the model (Blocks 1, 4, 6, and 10, considering that both ViT-S and ViT-B have a total of 12 blocks). In this case, the input to the binary classifier is the concatenated multi-head output of the selected blocks, which is directly processed by the fully connected layers without passing through a convolutional part (see Appendix for architecture details). As shown in Table III, the attack remains effective even when applied to transformer models, significantly reducing classification performance despite having access to only a limited portion of the model, which is assumed to be on the edge side.
Model Class Rand 1/4 2/4 3/4 4/4
ViT-S 0 71.6 40.24 27.30 19.28 12.67
109 - 36.52 37.34 25.39 14.61
343 - 47.78 41.90 29.67 23.32
ViT-B 0 76.1 55.40 52.95 44.60 32.41
109 - 53.13 47.47 30.81 16.21
343 - 56.68 55.72 51.57 42.12
TABLE III: Accuracy on ImageNet for ViT-S and ViT-B under edge-side attacks with $\epsilon=16/255$, considering different depths and target classes (’tench’-0, ’brain coral’-109, ’warthog’-343). For comparison, accuracy with respect to random $\epsilon$ noise is also reported.
Untargeted Attack (Accuracy $\downarrow$) Targeted Attack (Success Rate $\uparrow$)
Edge CosUAP (Depth) - Sec. V-C Black-Box Transfer (Attack Model) EdgeUAP (Depth) Sec. V-B Black-Box Transfer (Attack Model)
Model 1/4 2/4 3/4 4/4 MNV2 RN50 VGG16 WR110 1/4 2/4 3/4 4/4 MNV2 RN50 VGG16 WR110
MNV2 0.62 2.58 2.60 0.20 0.70 11.15 19.93 14.31 4.26 4.26 7.04 67.66 97.91 34.39 32.11 24.90
RN50 0.68 2.58 1.36 0.86 45.24 0.68 22.83 8.708 0.14 10.27 38.63 90.17 15.61 94.30 31.67 59.87
VGG16 0.82 1.52 0.82 0.48 28.76 8.76 0.44 11.47 0.06 1.78 44.62 80.36 66.11 41.32 96.90 28.17
WR110 20.98 10.33 4.72 2.32 63.28 20.04 34.91 1.80 0.16 7.78 40.94 74.51 6.08 36.75 10.89 89.22
TABLE IV: Comparison of Edge-Only and Black-Box models under untargeted and targeted attacks. Classification accuracy (↓) is reported for untargeted attacks, and Targeted Success Rate (TSR) (↑) for targeted attacks, targeting class 109 (brain coral). Red cells denote the optimal transferability given by a match between surrogate and target model (white-box setting).
Robust Model Clean Acc Threat Model Classes
1 2 3 4 5 6 7 8 9 10
ResNet50 83.28% Edge-1/4 54.6 54 50.8 50.6 57.6 56.9 40.1 64.3 69.1 57.2
Edge-2/4 31.1 46.1 38.7 43.2 43.6 51.7 21.3 50.1 21.3 48.6
Edge-3/4 24.9 35.8 35.3 30 33.6 36.7 14.7 34.6 40.1 44.4
Edge-4/4 15.8 27.6 26.3 28.6 22.9 30.2 10 24.3 35.3 39.1
UAP 18 34.6 25.3 23.3 21.7 31 18.8 32.9 32.4 34.8
MobileNetV2 83.35% Edge-1/4 45.2 31 58.9 48.8 35.2 59.1 30.1 42.8 46 22.1
Edge-2/4 24.8 22.4 36.8 35.9 27.3 37.8 14.8 29.8 23.2 19.9
Edge-3/4 23.6 18.2 32.8 30.6 22 36.8 15 24.4 22.8 10.2
Edge-4/4 18.8 19.6 31 30.4 22.2 31.4 14.1 21.9 19.4 7.4
UAP 26.3 20.8 27 32.8 22.6 31.7 20.9 27 23.1 14.9
VGG16 85% Edge-1/4 40 41.9 26.1 34.8 36 43.9 18.2 37.2 32.2 33.3
Edge-2/4 25.8 33.2 17.3 27.4 19.6 30.7 9.8 16.8 16.6 21.7
Edge-3/4 17.1 23.1 13.8 24.4 14.1 26.8 4.2 15.4 14.9 10
Edge-4/4 11.3 22.3 10.4 19.3 13 22.1 3.4 10.7 13.7 7.3
UAP 13.6 19.6 12 9.5 16.4 20 12.3 16 16.7 14.8
TABLE V: Evaluation of the effectiveness of adversarial training on CIFAR-10 against edge-only attacks performed with $\epsilon=16/255$. Note that we conducted targeted attacks against all 10 classes of CIFAR-10, represented in the plot with different colors, considering the various edge depths studied in previous experiments. The final bars in the plot indicate the performance of untargeted UAPs.
VI-G Threat Models Comparisons
To better understand the impact of the presented threat model, we provide a comparison with a black-box adversarial setting in Table IV, where, differently from our assumption, the attacker does not have access to the edge part of the target model but can leverage a locally available architecture that has been trained on a similar data distribution, hoping to achieve high transferability to the unknown target model. In the following, we distinguish the analysis of these comparisons between untargeted and targeted attacks.
Untargeted attack comparison. For the untargeted setting (left part in Table IV, we compare the two threat models by using, for the Edge-Only scenario, the CosUAP attack in the edge setting (Sec. V-C), computed across different edge depths. For the black-box scenario, we consider the UAP attack [9], testing different surrogate models (red cells indicate cases where the surrogate and target models are the same, i.e., the ideal white-box setting). From the results, CosUAP-Edge attacks are often more effective than black-box transferability, showing that, in split-inference scenarios where accessing the edge is possible, this provides a more reliable attack.
Targeted attack comparison. For the targeted analysis (right part in Table IV), extending the study in Fig. 4, we present a direct comparison of the targeted approach proposed in Sec. V-B with the transferability of UAPs computed in a targeted setting (the target class considered is “brain coral”). In this case, the benefits of EdgeUAP are less evident and strongly depend on the available depth. In black-box scenarios, a high targeted success rate is mainly due to the architectural similarity between the surrogate and target models (e.g., ResNet and WideResNet). However, targeted attacks in EdgeUAP can be comparable to strong black-box attacks when the depth of the edge model is sufficiently large (e.g., $3/4$ or $4/4$).
VI-H Analysis and discussion of defense mechanisms
Beyond demonstrating the feasibility and effectiveness of the proposed attacks, it is important to consider potential defense mechanisms and their limitations in the presented setting. Several works have proposed defenses for split and collaborative inference, including feature obfuscation, noise injection, pruning, and representation compression, mainly to mitigate privacy leakage and reconstruction or inversion attacks [53, 54, 55]. However, these defenses are typically not designed to address adversarial robustness. In particular, most existing approaches focus on protecting data confidentiality rather than preventing the purposeful manipulation of intermediate representations by evasion attacks. This highlights the need for novel defense strategies explicitly tailored to edge-accessible adversarial threat models.
A widely adopted mechanism for improving the robustness of generic neural networks against adversarial examples is adversarial training, which has also been extended to the domain of universal perturbations [56]. To evaluate the impact of our attack in this setting, we trained models from scratch on the CIFAR-10 training set using an adversarial training regime based on untargeted UAPs with perturbation magnitude $\epsilon=8/255$, aiming to balance clean accuracy and robustness to universal attacks [56]. We then evaluated the models on the CIFAR-10 test set and launched edge-only attacks at inference time using $\epsilon=16/255$. As shown in Table V, while adversarial training improves the overall robustness of the model, our edge-only attack, evaluated in the targeted setting across all CIFAR-10 classes, often gets a close effectiveness of white-box UAPs (e.g., for layers 2, 3, and 4 when targeting class 6), and in some cases even surpasses it. Ideally, a robust model should exhibit increased resistance in the early layers, so that adversarial effects do not propagate to deeper representations. However, results in Table V show that perturbations crafted using only edge-layer information can still strongly influence downstream predictions. These findings suggest that standard adversarial training could be further adapted and explored to address and mitigate edge-side adversarial threats.
VII Discussion and Conclusion
This work introduces a new threat model in the field of distributed learning and inference that enables edge-only adversarial attacks. In particular, we define two approaches for generating UAPs for both targeted and untargeted attacks without requiring any knowledge of the cloud-side model and the full output class distribution.